# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

### v1

@@ 3.2. Code Clone Detection
Buffer overflow strcpy, strncpy, memcpy, memset, read, write, gets, gets_s, strcat
@@
3.2.1. AST-level semantic comparison
As exemplified in example 2 of Section 2.3, Lines 5-6 (deleted pre-commit lines) and 7-8 (added line of post-commit, e.g., included in $S_{v}$) exhibit high semantic similarity. However, previous research failed to identify this high similarity through simple syntactic analysis, resulting in the commit (version) being erroneously labeled as an introducing commit (vulnerable version). We propose a semantic similarity comparison method at AST-level to address this limitation. AST is a tree representation of code that preserves well-defined components of statements, explicit statement order and the execution logic, which contains rich semantic information (Yang et al., 2021).
1) Function inlining. Due to code refactoring often integrating code snippets into a new function, as shown in Listing LABEL:list:false_reason2, the codes in Lines 7-9 are integrated into functions SSL2_RECORD_add_length, SSL2_RECORD_add_off and SSL2_RECORD_get_length. Therefore, we utilize the inline technique to expand functions during AST generation to gather more semantic information about the statements. This involves obtaining the function definition at the callsite and expanding the function body inline, resulting in a more comprehensive code representation. We get the AST for the target file using Clang without performing actual compilation. Furthermore, through recursive retrieval of line numbers for sub-ASTs, we can locate the sub-AST corresponding to $S_{v}$.
1@@ -567,18 +569,18 @@ int dtls1_read_bytes(SSL *s, int type, unsigned char *buf, int len, int peek)
2%diff --git a/ssl/record/rec_layer_d1.c b/ssl/record/rec_layer_d1.c
3%index 80f10be480222..c58af7f36abee 100644
4%--- a/ssl/record/rec_layer_d1.c
5%$+$ b/ssl/record/rec_layer_d1.c
6 if (!peek) {
7- rr->length -= n;
8- rr->off += n;
9- if (rr->length == 0) {
10+ SSL3_RECORD_add_length(rr, -n);
11+ SSL3_RECORD_add_off(rr, n);
12+ if (SSL3_RECORD_get_length(rr) == 0) {
13 s->rlayer.rstate = SSL_ST_READ_HEADER;
14
15%In ssl/record/record_locl.h
16#define SSL3_RECORD_add_length(r, l) ((r)->length += (l))
17#define SSL3_RECORD_add_off(r, o) ((r)->off += (o))
18#define SSL3_RECORD_get_length(r) ((r)->length)
Listing 3: Example of Code Outlining Modification.
2) AST normalization. In programming, different coders may have different coding styles, resulting in variations in code structure. To ensure consistency and minimize noise introduced by coding style differences, it is imperative to normalize the AST. AST normalization involves conditional structure normalization, loop structure normalization and relational operation normalization.
  • Loop structure can be represented as while, do-while and for structures, we normalize all the loop structure to while (Xue et al., 2022). The for structure is for(initialize; condition; increment) {statements;}, we transform it into while structure: initialize; while(condition) {statements; increment}.
  • Relational operation include commutative and noncommutative operation. Commutative relational operation (i.e., +, !=, &&) refers to the operation where commutate the operands does not change the operation semantics. We normalize the two operands (left and right child nodes) of the operation (parent nodes) by alphabetical order. For example, we transform $b+a$ (b is the left child node) to $a+b$ (a is the left child node). Noncommutative relational operation (i.e., $>$, <=<=< =) will change the operation semantic when commutating the operands, we normalize them by predefined rules. For example, we transform $b<a$ to $a>b$ (normalize operation $<$ to $>$).

### v2

@@ III-B3 Statement Weight Allocation
Buffer overflow
strcpy, strncpy, memcpy, memset, read,
write, gets, gets_s, strcat
@@
Parameters. There are two parameters for the weight allocation of statements, i.e., $W_{sensitive}$ and $W_{base}$, that could affect the effectiveness of Vercation: 1) $W_{base}$ is the base weight for initializing the weights of all statements. 2) $W_{sensitive}$ is the weight for the statements containing sensitive functions. We set $W_{sensitive}=2\times W_{base}$, $W_{base}=1.0$, the detailed weighting configuration experiment is in Section V-B.
III-B4 Code Clone Determination
We calculate the similarity_score for each code modification of the previous commit with the backtrack. We set a threshold $\vartheta_{3}$, if $similarity\_score<\vartheta_{3}$, we determine that there is no code clone between the pre-modification code of this previous commit and the vulnerable statements, thus confirming that this commit introduced the vulnerability. Otherwise, we proceed with the commit backtrack process.
$\begin{split}similarity\_score=\frac{sim\_a\times W_{sensitive}+sim\_b\times W_{base}}{a\times W_{sensitive}+b\times W_{base}}\end{split}$
where, $W_{sensitive}$ and $W_{base}$ are predefined weights assigned to different types of statements; $a$ and $b$ are the total counts of the statements in $S_{v}$ corresponding to these weight types; sim_asim\_a and sim_bsim\_b are the counts of statements that were found to be similar, either through syntactic or structural analysis.
III-C Vulnerable Version Range Determination
To precisely determine the range of software versions affected by a vulnerability, we employ a methodology based on the principle of commit reachability. The core premise is that a vulnerability, once introduced in a specific vic, is propagated to all subsequent commits in its lineage until it is fixed by a vfc.
The basis for our analysis is the concept of reachability, where a commit A is considered reachable from another commit B, if commit A is an ancestor of commit B in the project’s commit graph [17]. As version tags are essentially pointers to specific commits, they inherit the reachability properties of the commits they represent. This allows us to map the vulnerability’s presence from the commit level to software versions.
Our methodology formally defines the set of vulnerable versions ($V_{v}$) as follows: First, we identify $V_{i}$, the set of all version tags reachable from vic. Second, we identify $V_{f}$, the set of all version tags reachable from vfc. The vulnerable version set $V_{v}$ is then derived from the set difference between these two sets: $V_{v}=V_{i}-V_{f}$.
We illustrate the procedure by the case of CVE-2021-20294 in the Binutils project [51]. For this vulnerability, the set of versions containing the vic is [binutils-2_35, binutils-2_42], while the set of patched versions containing the vfc is [binutils-2_36, binutils-2_42], with binutils-2_42 being the latest version of the Binutils project. Therefore, we consider the version tags [binutils-2_35, binutils-2_36) as vulnerable. This notation signifies a range that includes binutils-2_35 but excludes binutils-2_36.
IV Experimental Setup
IV-A Dataset
With confirmation from the author of V-SZZ, it is established that V-SZZ assumes the deleted lines in the vfc as vulnerable codes and manually labels the first introduction of deleted lines in a vfc as the vic to construct the dataset. However, as mentioned in Section II-C, deleted lines cannot fully capture the vulnerability logic. Additionally, it cannot be guaranteed that the identified software version will necessarily exhibit the vulnerability. Meanwhile, V-SZZ excludes the vfcs that only contain added lines of code in their vulnerability selection, as the SZZ algorithm relies on tracking deleted lines of code to locate vic and thus cannot handle such cases. Therefore, we need to construct a ground-truth dataset with labeled vulnerable versions or non-vulnerable versions of OSS, which are verified by a public PoC.
Vulnerabilities. To build a reliable evaluation dataset, we formulate four vulnerability selection criteria: 1) The vulnerability must be a CVE reported in an OSS project; 2) The vulnerability must have publicly available PoCs and patches; 3) The PoC must be reproducible across different versions of the OSS and 4) The OSS must have a certain number of versions for comprehensive analysis. Based on the selection criteria, we established a dataset comprising 122 publicly disclosed CVEs associated with public PoC and patches, spanning 12 prevalent OSS, as detailed in Table II. The published date of these CVEs from 2016 to 2025. Meanwhile, the collected vulnerabilities cover 13 common CWE types. Some vulnerabilities are classified into multiple CWEs. For example, CVE-2023-1579 belongs to both CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer) and CWE-787 (Out-of-bounds Write). Among the dataset, vulnerabilities belonging to CWE-787 (Out-of-bounds Write) are the most numerous, accounting for 20.27%.
TABLE II: Ground-truth Dataset Overview.
IDX Name #Version #CVE Domain
1 Binutils 157 18 Programming tools
2 cJSON 48 3 JSON parser
3 FFmpeg 405 31 Multimedia processing
4 Jasper 99 12 Image coding toolkit
5 Libarchive 51 10 Streaming processing
6 Libcaca 7 3 Graphics library
7 Liblouis 69 6 Braille translator
8 Libming 14 9 Flash library
9 Libtiff 78 7 Image tools
10 Libxml2 233 12 XML toolkit
11 OpenJPEG 25 8 Image codec
12 Pcre2 25 3 Regular expression engine
Total - 1,211 122 -
Patches. We acquired security patches by crawling vfc from the OSS Git repositories. In total, the vfcs cover 2,287 modification lines containing 1,532 insertion lines and 755 deletion lines. Among these, 289 lines are unrelated to vulnerability (e.g., modifications in the ChangeLog file). On average, each vfc has 21.8 modification lines. The number of added lines significantly exceeds that of deleted lines, underscoring the critical importance of incorporating additional code into the analysis process. Among these, 10.3% are large vfcs (the number of modification lines exceeding 50). Notably, 79 vfcs (64.75%) contain both code insertion and deletion, 29 vfcs exclusively insertions (23.77%), and 15 vfcs solely involve deletions (12.30%).
  1. (i)
    The fixed version can be determined through the official release of vfcs. We confirm whether versions following the fixed version are vulnerable by checking if they contain the patch code.
  2. (ii)
    For the remaining version, we obtained the public PoC of each vulnerability from the Git issues or by referring to OSS sites.
  3. (iii)
    To validate the PoC on each corresponding OSS version, we establish the dependent environment for each OSS version and compile the libraries.
  4. (iv)
    During the PoC input testing process for each vulnerability, we analyze the triggering conditions and dangerous behaviors of the vulnerability and label the statements that trigger the vulnerability as $S_{trigger}$. For versions where the PoC cannot trigger the vulnerability (as some PoCs are only applicable to certain specific versions), we check whether the version contains $S_{trigger}$ to label the version.
In our experimental setup, we took 10 days to build the software and version pool. Then we took 30 days to build the ground-truth dataset, including the time taken to establish the dependent environment for each software version, as well as the time to perform vulnerability verification and labeling.
IV-B Baseline and Metrics
Baseline. For the vulnerable version identification task, we selected the NVD [52], along with SOTAs: SZZ-based methods (AG-SZZ [53], B-SZZ [54], V-SZZ [17]) and clone-based methods (V0Finder [16], V1SCAN [27], VERJava [28] and Vision [29]) for accuracy comparison. We gathered the vulnerable version range of each CVE from CPE to evaluate the NVD’s accuracy.
The basic idea of SZZ algorithms is to backtrack the commit history to locate the earliest commit that introduces the deletion lines removed by security patches. We use different SZZ algorithms to identify the vic, with the versions between the vic and the vfc considered vulnerable. Since these algorithms trace deleted lines individually, instances may arise where multiple vic are identified, and in such cases, we select the earliest one as the detection outcome.
V0Finder, V1SCAN, VERJava, and Vision are clone-based methods. We use their techniques to detect code clone for each version and identify the affected versions of the vulnerabilities in the ground-truth dataset. Because VERJava and Vision are applied in Java, we modified the code to make them compatible with C/C++.
For Vercation, we selected thresholds $\vartheta_{1}=0.9$, $\vartheta_{2}=0.8$ and $\vartheta_{3}=0.7$ (related experiments are introduced in Section V-C). In terms of the performance of vulnerability code extraction, we compared the differences between combining various LLMs (GPT-4, CodeLlama-13B, DeepSeek-V3) and using the static analysis tool Joern parser [33] alone.
Metrics. To evaluate the accuracy, we employ the following three metrics, i.e., true positives ($TP$), false positives ($FP$), false negatives ($FN$), precision ($\frac{\#TP}{\#TP+\#FP}$), recall ($\frac{\#TP}{\#TP+\#FN}$) and F1-score ($\frac{2*precision*recall}{precision+recall}$), to measure the accuracy of the above methodologies. These metrics are also used in previous studies to evaluate the technique’s performance [17, 9]. Because of significant differences in the version count for distinct OSS, we compute the precision and recall for each OSS and report the average of precisions and recalls finally.
IV-C Implementation Details
We utilized Joern [33] for program slicing and developed Python scripts to extract control and data dependencies, outputting the sliced statements as dangerous flows. We explored three LLM models, GPT-4 [30], CodeLlama-13B [31], and DeepSeek-V3 [32] for vulnerable code extraction. We use the public API developed by OpenAI to perform the experiment in GPT-4. The API version is GPT-4-0613 published in 2023. CodeLlama was created through further fine-tuning of Llama 2 on specific code datasets. We selected the 13B parameter model (codellama/CodeLlama-13b-hf) from the CodeLlama model series. The API of DeepSeek-V3 was published in July 2024. We set the temperature parameter to 1.0. To account for the stochastic nature of LLM outputs, we generate 10 independent responses for each CVE [55]. We employ a majority voting strategy where statements appearing in $\geq$ 6 out of 10 runs are included in the final vulnerable statement set for evaluation. The maximum size of tokens is 1,024. We deployed the Codellama model on our server with four NVIDIA RTX A5000 GPUs. Similar to GPT-4, the temperature parameter is set to 1.0, and the maximum size of tokens is set to 1,024. Subsequently, we apply weight allocation strategies to extract vulnerable statements using Python scripts. For threshold selection, we used $\theta_{1}$ = 0.9, $\theta_{2}$ = 0.8 and $\theta_{3}$ = 0.7 based on grid search optimization described in our sensitivity analysis. In the code change detection, we utilize the Clang 10.0.0 tool for function expansion (inline optimization) and AST generation of C/C++ code. Specifically, we use the -fsyntax-only option to perform syntax checking without actual compilation. We then develop Python scripts to achieve AST normalization and similarity comparison. Overall, we constructed our system with 6K LoC in Python.
V Evaluation
In this section, we evaluate the performance of the proposed Vercation by answering the following research questions (RQs).
  • RQ1. Overall Effectiveness: How does the overall performance of Vercation compare against state-of-the-art vulnerable version identification methods?
  • RQ2. Architectural Contribution: How do the key architectural choices in Vercation—the use of an LLM, static analysis, and statement weighting—contribute to its performance?
  • RQ3. Component Robustness: How accurate and robust is Vercation’s statement-level code clone detection?
  • RQ4. Real-World Application: What are the effects of applying Vercation in the real world scenarios?
V-A RQ1: Overall Effectiveness
We first compare Vercation against the NVD and SOTA methods. Table III shows the vulnerable version identification results of NVD, SZZ algorithms, V0Finder, V1SCAN, VERJava, Vision, and Vercation. Note that some CVEs do not provide vulnerable versions in NVD, such as CVE-2021-30499 [56]. SZZ algorithms cannot work on the vfcs containing solely added lines (i.e., 11 cases), so we exclude these cases in approaches using the SZZ algorithm.
V-A1 Comparison with NVD
The NVD achieves a precision of 66.8% and a recall of 41.8%. False positives often occur because NVD may broadly flag all versions preceding a patch as vulnerable without specific analysis. This situation exists in 18 CVEs (15%) of the ground truth. For instance, security experts found a buffer overflow risk exists in Libtiff 4.4.0, but NVD simply reported the versions of Libtiff<<4.4.0 were exposed in vulnerability. In reality, the vulnerable version range is Libtiff<<4.4.0 and $\geq$4.0.0.
The reason for low recall in NVD can be summarized into two types: (i) NVD only reported the version in which vulnerabilities were found. For example, CVE-2017-14152 [57] was found in OpenJPEG 2.2.0, then NVD reported the affected version was only 2.2.0, whereas the real vulnerable version range was 2.2.0-2.2.1. The problem exists in 57 CVEs (47%) of the ground truth. (ii) Some vulnerabilities are not fixed promptly after disclosure but are addressed after several versions have been released. These vulnerable versions can be overlooked if NVD does not update the version information. For example, CVE-2021-33815 [58] was discovered in FFmpeg 4.4 but was fixed in version 5.0. Due to the lack of timely tracking of the fix information, NVD overlooked the vulnerable versions 4.4.1-4.4.4.
TABLE III: Comparison with NVD Database and SOTA works.
Type Methods Precision Recall F1 score
- NVD Database 0.668 0.418 0.514
SZZ
Algorithm
AG-SZZ [59] 0.372 0.625 0.466
B-SZZ [54] 0.378 0.543 0.446
V-SZZ [17] 0.756 0.851 0.801
Clone-based
Approach
V0Finder [16] 0.829 0.745 0.785
V1SCAN [27] 0.863 0.724 0.788
VERJava [28] 0.891 0.247 0.386
Vision [29] 0.842 0.881 0.861
- Vercation 0.918 0.945 0.931
V-A2 Comparison with SZZ-based Methods
The basic idea of SZZ algorithms is to backtrack the commit history to locate the earliest commit that introduces the deletion lines removed by patches. Regarding F1-score, Vercation improves the best-performing baseline V-SZZ by 16.2%. And the difference between the F1 of B-SZZ and AG-SZZ is small (0.446 vs. 0.466). We affirmed that there are two main causes of false alarms in these approaches:
  1. (i)
    The line mapping algorithm based on edit distance does not consider the behavior of the source code. As shown in Listing LABEL:semantic, The deleted lines were encapsulated within the added functions, and in fact, they have the same behavior. Line mapping algorithms failed in these cases, unable to identify the true vic.
  2. (ii)
    The SZZ algorithm treats each deleted line as an independent origin for backtracking, resulting in multiple vic being identified. For example, the security patch of CVE-2020-35965 [60] from FFmpeg has 5 deleted lines. The true vulnerability logic lies in the lack of checking the size of ymax before executing the memset zero operation, potentially resulting in out-of-bounds writes to memory. Our method utilizes a weighted allocation approach to increase the weight of memset operation, thus backtracing to the correct vic. However, SZZ algorithms separately backtrack different deleted lines, resulting in the identification of 5 different vic, significantly reducing the precision.
V-A3 Comparison with Clone-based Methods
In the same dataset, V0Finder achieved a precision of 82.9% and a recall of 74.5%. There are two reasons for the inaccuracies of V0Finder: 1) V0Finder generates a hash value for the entire vulnerable function as a vulnerability fingerprint. This fingerprint is coarse because the vulnerable function contains a lot of vulnerability-irrelevant code, introducing a significant amount of noise. 2) V0Finder considers the function containing all deleted patch code as a vulnerable function clone. This approach also overlooked the inserted code of the security patch.
V1SCAN achieves an F1-score of 0.788, combining version-based detection with code-based detection methods. For version-based detection, V1SCAN obtains CVE-affected OSS versions from NVD’s CPE database and filters out functions belonging to versions outside the CVE’s affected range. As we mentioned earlier, there are many errors in NVD’s affected version information, which consequently affects V1SCAN’s detection results. For code-based detection, V1SCAN primarily relies on text similarity comparisons, overlooking deeper relationships between code snippets. When functions are refactored or outlined into new functions, this approach fails to identify vulnerability-related code, resulting in false negatives.
Table III reports a low recall of 0.247 in VERJava. VERJava has a fundamental issue in determining vulnerable versions: it requires the target version’s code to simultaneously meet two strict conditions - perfect matching of deleted lines in the vfc (delSim $\geq$ 1.0) and the absence of most added patch code (addSim $\leq$ 0.9). In project development, code frequently undergoes changes, including variable name modifications and function refactoring. This means that even versions that contain vulnerabilities are unlikely to meet such strict matching requirements due to code evolution. As a result, many vulnerable versions are incorrectly identified as safe versions, leading to serious false negatives. Another problem with VERJava is similar to V0Finder, it incorrectly assumes that deleted lines in the patch represent vulnerable features.
Vision achieves an F1 of 0.861, combining critical statement selection with weighted dependency graphs. Vision relies on taint analysis as a first step for critical statement selection. However, this approach often includes many statements that are control or data-dependent but not actually relevant to the vulnerability, which consequently affects Vision’s precision in extracting vulnerable code. Moreover, Vision’s analysis is limited to Java packages from Maven repositories, which may not fully capture the vulnerability patterns in other languages.
Accuracy of Vercation. Vercation demonstrates consistent performance across most projects, achieving high precision (>> 0.90) on 9 out of 12 projects and perfect recall (1.00) on several projects, including cJSON and Libcaca. Notably, while baseline methods like V-SZZ and V0Finder show significant performance fluctuations across different projects, Vercation maintains more stable performance. For instance, on the Binutils project, which has complex version management and frequent refactoring, Vercation achieves a 0.95 F1 score, outperforming V-SZZ (0.47 F1 score) and V0Finder (0.88 F1 score).
Though Vercation performed well in the version identification, it still failed to identify inducing commits for a small number of vulnerabilities. One reason is that in some cases, slight modifications in the AST have little impact on the similarity between the AST before and after the modification, but are crucial in introducing vulnerabilities. Another reason is that despite using the weight allocation strategies, there are still some statements related to patch variables that are not relevant to the vulnerability but have been assigned high weights. The above two reasons may incorrectly identify vic, leading to a misidentification of the vulnerable version range.
In our evaluation, 5.7% of traced commit histories involved merge commits with multiple parents. For example, in CVE-2022-1355 [61] from Libtiff, the vulnerability was introduced in a feature branch that was later merged into the main development line, creating a non-linear history with two parent commits at the merge point. For these cases involving non-linear histories, Vercation achieved a precision of 91.2% and recall of 92.5%, demonstrating that our multi-path traversal strategy effectively handles branching scenarios without significant performance degradation.
RQ1: Vercation demonstrates superior performance in vulnerable version identification on our ground-truth dataset, achieving an F1 score of 93.1%. This significantly outperforms SOTA works. The prompt strategy combining Few-shot and CoT techniques performs the best.
V-B RQ2: Architectural Contribution
V-B1 Program Slicing & LLM Ablation Study
We conducted ablation experiments by separately removing the program slicing component (implemented by Joern) and the LLM component (DeepSeek) from phase 1.
(a) Extracted by Using Only Joern Parser.
(b) Extracted by Using Only LLM.
(c) Extracted by Combining Joern Parser and LLM.
Figure 5: The Vulnerable Statements Extracted by Different Methods.
TABLE IV: Performance of LLMs in Vulnerability Logic Comprehension.
Method (Tool) Avg #Vuln Stmts1 Precision Recall F1
Static Analysis (Joern)
22.64 0.435 0.544 0.483
LLM (DeepSeek)
11.33 0.742 0.796 0.768
Static Analysis + LLM
(Joern + DeepSeek)
5.66 0.918 0.945 0.931